4.Under good weather conditions, 80% of flights arrive on time. During bad weather, only 30% of flights arrive on time. Tomorrow, the chance of good weather is 60%. What is the probability that your flight will arrive on time?
Answer by robertb(5830)

P(on time|good weather) = 0.80
P(on time|bad weather) = 0.30
P(good weather) = 0.60 ==> P(bad weather) = 0.40
Then
P(on time) = P(on time AND good weather) + P(on time AND bad weather)
=P(on time|good weather)*P(good weather) + P(on time|bad weather) * P(bad weather)
=0.80*0.60 + 0.30*0.40 = 0.60, the probability that your flight will arrive on time.